OpenStudy (anonymous):

What is the solution of -2x + 8 = 20?

OpenStudy (anonymous):

well solve for x.. -2x + 8 = 20 get all the numbers without variables on one side -2x = 20 - 8 -2x = 12 then divide both sides by -2 to get the variable by itself and you'll get x = -6

OpenStudy (anonymous):

what is the solution of x+5/2=15

OpenStudy (anonymous):

do the same thing, subtract 5/2 from both sides and there's your answer. all you have to do is get x by itself

OpenStudy (anonymous):

then that gives you x=5/2*15?

OpenStudy (anonymous):

you don't want to multiply, subtract

OpenStudy (anonymous):

15-(5/2) = 25/2

OpenStudy (anonymous):

that would only be the case if you it was 2/5x = 15

OpenStudy (anonymous):

What is the solution of -(x + 1) - 5 > 10? okay and how would i work this problem out?

OpenStudy (anonymous):

first distribute the negative

OpenStudy (anonymous):

then combine like terms, and do whatever you need to do to get x by itself

OpenStudy (anonymous):

which would give me x-4>10?

OpenStudy (anonymous):

just like the other problems. if you have x plus something, subtract it from both sides, if you have something times x then divide both sides by it. if you have something times x plus/minus something, add/subtract from both sides than divide

OpenStudy (anonymous):

no..when you distribute the negative be very careful -(x+1)= -x-1

OpenStudy (anonymous):

then -x-1-5 = -x-6

OpenStudy (anonymous):

then that the answer?

OpenStudy (anonymous):

no. get x by itself. your subtracting 6 here so you want to add x to both sides and then divide by -1. try focusing on the method instead of the answer. that's how you learn

OpenStudy (anonymous):

What is the quadratic function that is created with roots -10 and -4 and a vertex at (-7, -9)?